OpenStudy (anonymous):

how would you solve a proportion of 1/49w=w/9

OpenStudy (lgbasallote):

\[\huge \frac{1}{49w} = \frac w9?\]

OpenStudy (anonymous):

correct

OpenStudy (lgbasallote):

cross multiply \[\huge \frac{9}{49} = w \times w\] do you get this step?

OpenStudy (anonymous):

\[w=\sqrt{9/49}\]

OpenStudy (anonymous):

ah I did cross multiply but the exponent of w threw me off

OpenStudy (lgbasallote):

hmm did you do \[\huge w^2 = \frac{9}{49}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

vladislav mentioned \[\sqrt{9/49}\]

OpenStudy (anonymous):

this makes sense :D

OpenStudy (lgbasallote):

yup that's the next step..taking the square root of both sides \[w = \sqrt{\frac{9}{49}}\] \[w = \frac{\sqrt 9}{\sqrt{49}}\] i suppose you know the answer now?

OpenStudy (anonymous):

yes thank you very much ^_^

OpenStudy (lgbasallote):

you're welcome :)